# Unintended Impacts of LLM Alignment on Global Representation

### v1

@@ Abstract
Before being deployed for user-facing applications, developers align Large Language Models (LLMs) to user preferences through a variety of procedures, such as Reinforcement Learning From Human Feedback (RLHF) and Direct Preference Optimization (DPO). Current evaluations of these procedures focus on benchmarks of instruction following, reasoning, and truthfulness. However, human preferences are not universal, and aligning to specific preference sets may have unintended effects. We explore how alignment impacts performance along three axes of global representation: English dialects, multilingualism, and opinions from and about countries worldwide. Our results show that current alignment procedures create disparities between English dialects and global opinions. We find alignment improves capabilities in several languages. We conclude by discussing design decisions that led to these unintended impacts and recommendations for more equitable preference tuning.
@@ 1 Introduction
Recently, LLM-powered chat assistants OpenAI (2023a); Touvron et al. (2023); Tunstall et al. (2023b) have exploded in popularity. As of December 2023, ChatGPT has amassed over 100M weekly users OpenAI (2023b) and Llama-Chat-7B is downloaded almost one million times a month from HuggingFace111Llama-Chat-7B Huggingface Page. The success of these chat models is dependent on "alignment", which takes a base model with a language modeling objective and produces an instruction following preference-guided model to better serve user interests. Practitioners use algorithms such as RLHF Ouyang et al. (2022) and DPO Rafailov et al. (2023) to optimize models for attributes such as helpfulness and harmlessness and give them their chat assistant persona Ouyang et al. (2022); Bai et al. (2022); Zhu et al. (2023).
@@ 2 Related Work
Lambert et al. (2023) provides a fantastic overview of the risks of RLHF. Prior work has noted the social impacts of RLHF and Preference Tuning Liu (2023). Ouyang et al. (2022) identifies that they aligned their model with mostly English speakers from the US and Southeast Asia. RLHF has been observed to steer models towards outputs that are longer Singhal et al. (2023), more assertive Hosking et al. (2023), and less novel Kirk et al. (2023). RLHF can also make mistakes in the output more subtle Bai et al. (2022). Shaikh et al. (2023) finds that RLHF decreases grounding acts. Perez et al. (2023) find that RLHF makes models echo user opinions, stronger political views, and requests not to be shut down.
@@
| Tulu 2 DPO | DPO | GPT-4 | UltraFeedback | Tulu 2 | Mixed | Llama 2 | Internet Dump* |
| Starling LM | PPO (RLAIF) | GPT-4 | Nectar | OpenChat 3.5 | Mixed | Mistral v0.1 | Internet Dump* |
@@ 3 Alignment Process
In the supervised fine-tuning stage, the model is provided with prompts and example completions and fine-tuned to produce these sorts of completions. Popular SFT datasets for chat models include the human-written Flan222Note that Flan contains templated completions of other datasets rather than being fully naturally written Wei et al. and Open Assistant Köpf et al. (2023) datasets, and the synthetic ShareGPT 333https://sharegpt.com, Alpaca Taori et al. (2023), and Open-Orca Lian et al. (2023) datasets. All are variants of instruction following completions to task-oriented prompts. Typically, this step is used to make language models follow instructions rather than continue the input text based on the language modeling objective.
@@
After alignment, language models are either deployed inside a product or released for broader use. Notably, these models are a core technology that enables higher-level user-facing systems. While model developers may intend a specific audience, open-access models can be adopted anywhere and major LLM APIs are globally accessible444OpenAI and Google Supported Countries. As a result, due to the broad nature of their possible utility, even unintended impacts of alignment can affect their global adoption.
@@ 4 Global Representation: English Dialects
We report the different accuracies of LM guesses in Figure 3 with 95% confidence intervals. Whenever changes to performance are significant (p<0.05), the alignment steps increase US English performance much more significantly than other global Englishes. Before alignment, all Base models performed relatively the same across dialects (about 5% accuracy for Llama and 8% accuracy for Mistral). Though SFT improves performance across all dialects it creates a disparity in performance gains between dialects. For Mistral SFT, the difference between USA English and Indian English grows from 0.98% in Mistral to 15.0% in Mistral SFT, and for USA English and Nigerian, the difference grows from 1.3% to 10.3%. Similarly, for OpenChat, the disparity in USA English and Indian English grows from 0.98% to 16.9% and between USA and Nigerian English from 1.3% to 9.7%.
@@ 6 Global Representation: Opinions
We measure the probability of responding with each answer choice and compare the probability distribution with global respondents. Following the GlobalOpinionsQA paper Durmus et al. (2023) we measure 1-Jensen Shannon divergence between the LLM responses and responses for each country. We use a similar task setting to the original Anthropic paper. However, our analysis covers nine open models and all alignment stages, while the original analysis is limited to the Claude model. We report the change in similarity to global values in Figure 5 with 95% confidence intervals.
@@ 6.1 Reward Model Probing
Reward models are not well suited for multiple-choice assessments due to the limited response length. We build a dataset suited to counterfactual reward probing by collecting a set of 554 country-specific questions from the subreddit r/AskReddit 555https://www.reddit.com/r/AskReddit/. We search for questions using the queries "Which Country", "What Country", "Best Country", and "Worst Country" to collect varied questions. This resulted in 957 questions. After removing duplicates, questions with strictly factual answers, and questions that could not be answered with a specific country name, we were left with 554 quality-assured questions.
@@
We measure correlation with rankings by US citizens collected from Gallup polls in 2017 and 2023 Brenan (2023). Gallup surveyed 1,035 US adults in 2017 and 1,008 US adults in 2023 and asked them to rate countries as "Very Favorable," "Mostly Favorable," "Mostly Unfavorable," "Very Unfavorable," or "No opinion." The aggregate scores are used to compute a ranking over the 21 countries surveyed. We report the top 5 and bottom 5 countries from this list ranked by Starling in Table 3. Comparing just the rankings of these 21 countries to those produced by the Starling 7B RM, we find a 0.926 Spearman correlation with the 2017 results and a 0.849 Spearman correlation with the 2023 results. This indicates a high overlap between US opinions and the learned preferences of the Starling RM. These results offer a step towards answering the question, "To whose preferences are we aligning language models?" Western preferences certainly have a significant influence. We report all rankings by Starling RM along with a choropleth visualization in Appendix E. Unrestricted to the Gallup list, Starling ranks "Morocco," "the USA," "Slovenia," and "New Zealand" highest and "Western Sahara," "North Korea," "Turkmenistan," and "Central African Republic" lowest. Similar to our motivating example, "Where are you from?" we find the Starling model assigns low rewards to countries in central Africa and the Middle East.
@@ Acknowledgements
The authors would like to thank Omar Shaikh, Jared Moore, Vyoma Raman, Ananjan Nandi, Yanzhe Zhang, Matthias Gerstgrassar, Jing Huang, Yunze Xiao, Banghua Zhu, Chenglei Si, Daniel Campos, Rose Wang, and SALT Lab their feedback and suggestions at various stages of the project.

### v2

@@ Abstract
Before being deployed for user-facing applications, developers align Large Language Models (LLMs) to user preferences through a variety of procedures, such as Reinforcement Learning From Human Feedback (RLHF) and Direct Preference Optimization (DPO). Current evaluations of these procedures focus on benchmarks of instruction following, reasoning, and truthfulness. However, human preferences are not universal, and aligning to specific preference sets may have unintended effects. We explore how alignment impacts performance along three axes of global representation: English dialects, multilingualism, and opinions from and about countries worldwide. Our results show that current alignment procedures create disparities between English dialects and global opinions. We find alignment improves capabilities in several languages. We conclude by discussing design decisions that led to these unintended impacts and recommendations for more equitable preference tuning. We make our code and data publicly available on Github111https://github.com/SALT-NLP/unintended-impacts-of-alignment.
@@ 1 Introduction
Recently, LLM-powered chat assistants OpenAI (2023a); Touvron et al. (2023); Tunstall et al. (2023b) have exploded in popularity. As of December 2023, ChatGPT has amassed over 100M weekly users OpenAI (2023b) and Llama-2-Chat-7B is downloaded almost one million times a month from HuggingFace222Llama-2-Chat-7B Huggingface Page. The success of these chat models is dependent on "alignment", which takes a base model with a language modeling objective and produces an instruction following preference-guided model to better serve user interests. Practitioners use algorithms such as RLHF Ouyang et al. (2022) and DPO Rafailov et al. (2023) to optimize models for attributes such as helpfulness and harmlessness and give them their chat assistant persona Ouyang et al. (2022); Bai et al. (2022); Zhu et al. (2023).
@@ 2 Related Work
Lambert et al. (2023) provides a fantastic overview of the risks of RLHF. Prior work has noted the social impacts of RLHF and Preference Tuning Liu (2023). Ouyang et al. (2022) identifies that they aligned their model with mostly English speakers from the US and Southeast Asia. RLHF has been observed to steer models towards outputs that are longer Singhal et al. (2023), more assertive Hosking et al. (2023), and less novel Kirk et al. (2023). RLHF can also make mistakes in the output more subtle Bai et al. (2022). Shaikh et al. (2023) finds that RLHF decreases grounding acts. Perez et al. (2023) find that RLHF makes models echo user opinions, stronger political views, and requests not to be shut down. (Lin et al., 2024) explore approaches to mitigating the ’Alignment Tax’ where models that undergo RLHF lose skills they previously learned.
@@
| Tulu 2 DPO | DPO | GPT-4 | UltraFeedback | Tulu 2 | Mixed † | Llama 2 | Internet Dump* |
| Starling LM | PPO (RLAIF) | GPT-4 | Nectar | OpenChat 3.5 | Mixed ♠ | Mistral v0.1 | Internet Dump* |
@@ 3 Alignment Process
In the supervised fine-tuning stage, the model is provided with prompts and example completions and fine-tuned to produce these sorts of completions. Popular SFT datasets for chat models include the human-written Flan333Note that Flan contains templated completions of other datasets rather than being fully naturally written Wei et al. and Open Assistant Köpf et al. (2023) datasets, and the synthetic ShareGPT 444https://sharegpt.com, Alpaca Taori et al. (2023), and Open-Orca Lian et al. (2023) datasets. All are variants of instruction following completions to task-oriented prompts. Typically, this step is used to make language models follow instructions rather than continue the input text based on the language modeling objective.
@@
After alignment, language models are either deployed inside a product or released for broader use. Notably, these models are a core technology that enables higher-level user-facing systems. While model developers may intend a specific audience, open-access models can be adopted anywhere and major LLM APIs are globally accessible555OpenAI and Google Supported Countries. As a result, due to the broad nature of their possible utility, even unintended impacts of alignment can affect their global adoption.
@@ 4 Global Representation: English Dialects
We report the different accuracies of LM guesses in Figure 3 with 95% confidence intervals. Whenever changes to performance are significant (p<0.05), the alignment steps increase US English performance much more significantly than other global Englishes. Before alignment, all Base models performed relatively the same across dialects (about 5% accuracy for Llama and 8% accuracy for Mistral).
Though SFT improves performance across all dialects, it creates a disparity in performance gains between dialects. For Mistral to Mistral SFT, Indian English accuracy increased by 15.2%, Nigerian English accuracy increased by 20.3% and American English accuracy increased by 29.3%. Similarly, for Mistral to OpenChat, performance increased by 20.3%, 27.9%, and 36.3% for Indian, Nigerian, and American English respectively.
@@ 6 Global Representation: Opinions
We measure the probability of responding with each answer choice and compare the probability distribution with global respondents. Following the GlobalOpinionsQA paper Durmus et al. (2023) we measure 1-Jensen Shannon divergence between the LLM responses and responses for each country to measure agreement. We use a similar task setting to the original paper. However, our analysis covers nine open models and all alignment stages, while the original analysis is limited to the Claude model. We subtract LLM agreement with the USA from agreement with other nations and report the change in Figure 5 with 95% confidence intervals.
@@ 6.1 Reward Model Probing
Reward models are not well suited for multiple-choice assessments due to the limited response length. We build a dataset suited to counterfactual reward probing by collecting a set of 554 country-specific questions from the subreddit r/AskReddit 666https://www.reddit.com/r/AskReddit/. We search for questions using the queries "Which Country", "What Country", "Best Country", and "Worst Country" to collect varied questions. This resulted in 957 questions. After removing duplicates, questions with strictly factual answers, and questions that could not be answered with a specific country name, we were left with 554 quality-assured questions.
@@
We measure correlation with rankings by US citizens collected from Gallup polls in 2017 and 2023 Brenan (2023). Gallup surveyed 1,035 US adults in 2017 and 1,008 US adults in 2023 and asked them to rate countries as "Very Favorable," "Mostly Favorable," "Mostly Unfavorable," "Very Unfavorable," or "No opinion." The aggregate scores are used to compute a ranking over the 21 countries surveyed. We report the top 5 and bottom 5 countries from this list ranked by Starling in Table 3. Comparing just the rankings of these 21 countries to those produced by the Starling 7B RM, we find a 0.926 Spearman correlation with the 2017 results (p=1.78E-9) and a 0.849 Spearman correlation with the 2023 results (p=1.12E-6). This indicates a high overlap between US opinions and the learned preferences of the Starling RM. These results offer a step towards answering the question, "To whose preferences are we aligning language models?" Western preferences certainly have a significant influence. We report all rankings by Starling RM along with a choropleth visualization in Appendix E. Unrestricted to the Gallup list, Starling ranks "Morocco," "the USA," "Slovenia," and "New Zealand" highest and "Western Sahara," "North Korea," "Turkmenistan," and "Central African Republic" lowest. Similar to our motivating example, "Where are you from?" we find the Starling model assigns low rewards to countries in central Africa and the Middle East.
@@ Acknowledgements
This work was funded in part by a Meta grant and an NSF grant IIS-2247357. The authors would like to thank Omar Shaikh, Jared Moore, Vyoma Raman, Ananjan Nandi, Yanzhe Zhang, Matthias Gerstgrassar, Jing Huang, Yunze Xiao, Banghua Zhu, Chenglei Si, Daniel Campos, Rose Wang, and SALT Lab their feedback and suggestions at various stages of the project. Figure 2 has been designed using images from Flaticon.com.
@@ Appendix E Ask Reddit Full Results
Appendix F Additional Models
We validate our findings by experimenting with additional models. In particular, we experiment with Qwen1.5-7B versus Qwen1.5-7B-Chat Bai et al. (2023) and Yi-6B versus Yi-6B-Chat AI et al. (2024). We chose these models in particular as Western researchers did not align them, but rather Qwen was developed by Alibaba, and Yi was developed by 01.AI, two Chinese corporations. The Qwen model family does not release an intermediate SFT model only the final preference tuned model (RLHF with PPO), and Yi only performs SFT. We show results before and after alignment with the available models for each. We report results on all the same experiments as in the main paper.
F.1 Dialect
We report results for Qwen and Yi on the MD3 Dialect Intent Prediction Task in Figure 8. Similar to the results with the Western-aligned models, we find that the improvements are most significant for Yi for Indian English and American English: 7.41% and 3.98%, respectively, with only a 2.39% increase in Nigerian English. We see a drop in Qwen Chat’s performance. Upon qualitative inspection, the Qwen Chat Model outputs a preamble explanation for each guess, which causes it to exceed the 10 maximum new tokens. In order to account for this we run a special test setting for Qwen-Chat with it’s chat template and an increased max token limit of 100 tokens. Results of this run are presented in 9. These trends are more in line with the other models which tend to see improvements on the intent prediction accuracy in the aligned models. Similarly the best performance is seen in American English over the other dialects.
F.2 Languages
We report results for Belebele Reading Comprehension and TyDiQA Question Answering tasks in Figure 10. Qwen improves in reading comprehension in nearly all languages besides Telugu. Yi does not see significant increases in any language. The Qwen tech report evaluates multilingual benchmarks Bai et al. (2023) while Yi does not. Yi trains on the Chinese Open Instruction Generalist (COIG) dataset which is comprised almost exclusively of Chinese text Zhang et al. (2023). Although Qwen does not release its SFT data it is likely more multilingual data was included in Qwen’s SFT process.
F.3 Opinions
Finally, we run the GlobalOpinionsQA survey on both Qwen and Yi before and after alignment. We report the difference with USA alignment in Figure 11. We find similar results to the models assessed in the main text. Alignment to Jordan, China, and Nigeria, compared to the USA, typically remains more distant and decreases in the case of Yi. This contrasts with Germany and Australia, which have similar Jensen-Shannon similarity to the model compared to the USA. This finding is additionally interesting considering these models were aligned by researchers in China. It is worth noting that Yi did not undergo preference tuning, while Qwen underwent RLHF using PPO.